# A Free-fermion Formulation of Two-dimensional Ising Models

### v1

@@ III.1.2 Imaginary field
$~{}~{}~{}~{}\ln\left\{{\cosh}^{3}\left({2\beta J}\right)-1\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}\left.+{{\sinh}^{2}}\left({2\beta J}\right)\left[{\cos\theta+\cos\phi+\cos\left({\theta-\phi}\right)}\right]\right\}.$ (28)
One can examine that our result recovers the proposed solutions of this case [16, 22].
III.2 The triangular lattice
To deal with the triangular lattice Ising model, we employ the star-triangle transformation as shown in Fig. 5(a). The transformed spin system is on the honeycomb lattice, with the replacement of the nearest-neighbour interaction $J\to J^{\prime}$. $J^{\prime}$ is determined by Eq. (6). Applying the decorated lattice technique in the same way as that in Fig. 4(a), we construct a sixteen-vertex model of which the vertex site is shown Fig. 5(b). The mapping from the spin states to the arrow configurations is still that in Eq. (13). The relation between $J^{\prime}$ and $J^{\prime}_{1}$ is the same with that in Eq. (3).
III.2.1 Zero field
In the zero field, the partition function is simply
$Z={B^{N}}Z\left({\rm{honeycomb}},J\to J^{\prime},H_{\rm{ex}}=0\right)$
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z=\ln B+2\times{\rm{Eq}}.~{}{\rm{}}(21)\left({J\to J^{\prime}}\right).$ (29)
@@ III.2.1 Zero field
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z=\frac{1}{{8{\pi^{2}}}}\int_{0}^{2\pi}d\theta\int_{0}^{2\pi}d\phi\ln\left\{{e^{-6\beta J}}+3{e^{2\beta J}}\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}\left.-4\sinh\left({2\beta J}\right)\left[{\cos\theta+\cos\phi+\cos\left({\theta-\phi}\right)}\right]\right\}.$ (30)
This expression agrees with the known results of this case [32, 33, 34, 35].
III.2.2 Imaginary field
@@
$Z=i^{N}B^{N}A^{\prime 2N}\sum\limits_{\left\{{\tilde{s}}\right\}=\pm 1}{\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)}~{},$ (31)
with $A^{\prime}=\frac{1}{2}{e^{\beta J^{\prime}}}$ and
$\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)=\sum\limits_{{s_{1}}=\pm 1,{s_{2}}=\pm 1}(s_{1})e^{-\beta J^{\prime}{s_{1}}{s_{2}}}\times$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}{{e^{-\beta J^{\prime}_{1}\left[{{s_{1}}\left({{{\tilde{s}}_{1}}+{{\tilde{s}}_{2}}}\right)+{s_{2}}\left({{{\tilde{s}}_{3}}+{{\tilde{s}}_{4}}}\right)}\right]}}}.$ (32)
@@
$\omega_{1}=4{e^{-3\beta J^{\prime}}}{\left({{e^{-4\beta J^{\prime}}}-1}\right)^{{1\mathord{\left/{\vphantom{12}}\right.\kern-1.2pt}2}}},$
@@
$\omega_{3}=\omega_{4}=\omega_{5}=\omega_{6}=0,$
$\omega_{7}=-4{e^{-\beta J^{\prime}}}{\left({{e^{-4\beta J^{\prime}}}-1}\right)^{{1\mathord{\left/{\vphantom{12}}\right.\kern-1.2pt}2}}},$
@@
$\omega_{9}=\omega_{13}=2\left({{e^{-\beta J^{\prime}}}-{e^{\beta J^{\prime}}}}\right){\left({{e^{-4\beta J^{\prime}}}-1}\right)^{{1\mathord{\left/{\vphantom{12}}\right.\kern-1.2pt}2}}},$
$\omega_{10}=\omega_{14}=-\omega_{9},$
$\omega_{11}=\omega_{15}=-2\left({{e^{-\beta J^{\prime}}}+{e^{\beta J^{\prime}}}}\right){\left({{e^{-4\beta J^{\prime}}}-1}\right)^{{1\mathord{\left/{\vphantom{12}}\right.\kern-1.2pt}2}}},$
$\omega_{12}=\omega_{16}=-\omega_{11}.$ (33)
@@
$\tilde{\omega}_{9}=-2{e^{\beta J^{\prime}}}{\left({{e^{-2\beta J^{\prime}}}+1}\right)^{2}}{\left({{e^{-4\beta J^{\prime}}}-1}\right)^{{1\mathord{\left/{\vphantom{12}}\right.\kern-1.2pt}2}}},$
$\tilde{\omega}_{10}=-2{e^{\beta J^{\prime}}}{\left({{e^{-4\beta J^{\prime}}}-1}\right)^{{3\mathord{\left/{\vphantom{32}}\right.\kern-1.2pt}2}}},$
$\tilde{\omega}_{11}=-2{e^{\beta J^{\prime}}}{\left({{e^{-2\beta J^{\prime}}}-1}\right)^{2}}{\left({{e^{-4\beta J^{\prime}}}-1}\right)^{{1\mathord{\left/{\vphantom{12}}\right.\kern-1.2pt}2}}},$
$\tilde{\omega}_{12}=\tilde{\omega}_{14}=\tilde{\omega}_{16}=\tilde{\omega}_{10},$
$\tilde{\omega}_{13}=\tilde{\omega}_{9},$
$\tilde{\omega}_{15}=\tilde{\omega}_{11}.$ (34)
It is obvious that the odd free-fermion condition Eq. (12) is satisfied. Therefore, the triangular lattice Ising model in the imaginary field is equivalent with an odd free-fermion model. Now the result is
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z=i\frac{\pi}{2}+\ln B+2\ln A^{\prime}+\mathop{\lim}\limits_{{N_{{\rm{v}}}}\to\infty}\frac{1}{{{N_{{\rm{v}}}}}}\ln{Z_{{\rm{odd}}}},$ (35)
where the exact solution of the odd free-fermion model is given by Eq. (21) of Ref. [43]
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.+c\cos\phi+d\cos\left({\theta-\phi}\right)+e\left({\theta+\phi}\right)\right]$ (36)
@@
We should remark that the relation between our notation of weights and that of Ref. [43] is
@@
Substituting $A^{\prime}=\frac{1}{2}e^{\beta J^{\prime}}$ and Eqs. (6), (7) and (34) into Eq. (35), we give the final expression
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z=i\frac{\pi}{2}+\frac{1}{{16{\pi^{2}}}}\int_{0}^{2\pi}d\theta\int_{0}^{2\pi}d\phi\ln\left\{e^{-12\beta J}-4\right.$
$~{}~{}~{}~{}\left.+3e^{4\beta J}-8\sinh^{2}\left({2\beta J}\right)\left[{\cos\theta-\cos\phi+\cos\left({\theta-\phi}\right)}\right]\right\}.$ (38)
Then we have rederived the exact solution of this case [16, 22], and indicated that the model in this case can be seen as an odd free-fermion model.
It is easy to find that this frustrated system (when $J>0$) in the imaginary field has a finite residual entropy in the zero temperature limit
$S$ $=$ $\frac{1}{{16{\pi^{2}}}}\int_{0}^{2\pi}d\theta\int_{0}^{2\pi}d\phi\ln\left\{3-2\left[\cos\theta-\cos\phi\right.\right.$ (39)
@@
Here we omit the Boltzmann constant $k_{B}$. Interestingly, the residual entropy in the imaginary field is half that in the zero field [32] (we can also obtain the residual entropy in the zero field by Eq. (30), and compare the value with Eq. (39)).
@@ III.3.1 Zero field
$Z=\sum\limits_{\left\{{\left({{s_{1}},{s_{2}},{s_{3}},{s_{4}}}\right)}\right\}=\pm 1}{\omega\left({{s_{1}},{s_{2}},{s_{3}},{s_{4}}}\right)}$ (40)
@@
$\omega\left({{s_{1}},{s_{2}},{s_{3}},{s_{4}}}\right)=\sum\limits_{{s_{5}}=\pm 1}{{e^{-\beta J\left[{{s_{5}}\left({{s_{1}}+{s_{2}}+{s_{3}}+{s_{4}}}\right)+{s_{1}}{s_{2}}+{s_{3}}{s_{4}}}\right]}}}.$ (41)
@@
$\omega_{1}=\omega_{2}=2e^{-2\beta J}\cosh\left(4\beta J\right),$
$\omega_{3}=\omega_{4}=\omega_{5}=\omega_{6}=2e^{2\beta J},$
$\omega_{7}=\omega_{8}=2e^{-2\beta J},$
$\omega_{9}=\cdots=\omega_{16}=2\cosh\left(2\beta J\right).$ (42)
@@
$\tilde{\omega}_{1}=\frac{1}{2}\left(e^{-3\beta J}+3e^{\beta J}\right)^{2},$
$\tilde{\omega}_{2}=\tilde{\omega}_{3}=\tilde{\omega}_{4}=\tilde{\omega}_{5}=\tilde{\omega}_{6}=\frac{1}{2}\left(e^{-3\beta J}-e^{\beta J}\right)^{2},$
$\tilde{\omega}_{7}=\tilde{\omega}_{8}=\frac{1}{2}\left(e^{-3\beta J}+3e^{\beta J}\right)\left(e^{-3\beta J}-e^{\beta J}\right).$ (43)
One can see immediately that this case is an even free-fermion model. Making use of the even free-fermion solution as expressed in Eqs. (18) and (19), we give the exact result
@@
$=$ $\frac{1}{24{\pi^{2}}}\int_{0}^{2\pi}d\theta\int_{0}^{2\pi}d\phi\ln\left\{e^{-12\beta J}+18e^{-4\beta J}+24\right.$
$+21e^{4\beta J}+4\left(e^{-8\beta J}-e^{-4\beta J}-1+e^{4\beta J}\right)\left[\cos\theta\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.+\cos\phi-\cos\left(\theta-\phi\right)\right]\right\}.$
This result agrees with the known expressions (see Eq. (37) of Ref. [38] for instance).
III.3.2 Imaginary field

### v2

@@ III.2.1 Zero-field case
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z=\ln 2+\frac{1}{{8{\pi^{2}}}}\int_{0}^{2\pi}d\theta\int_{0}^{2\pi}d\phi$
$\ln\left[\cosh\left(2\beta J_{1}\right)\cosh\left(2\beta J_{2}\right)\cosh\left(2\beta J_{3}\right)\right$.
$~-\sinh\left(2\beta J_{1}\right)\sinh\left(2\beta J_{2}\right)\sinh\left(2\beta J_{3}\right)-\sinh\left(2\beta J_{3}\right)\cos\theta$
$~\left.-\sinh\left(2\beta J_{2}\right)\cos\phi+\sinh\left(2\beta J_{1}\right)\cos\left(\theta-\phi\right)\right].$ (46)
This expression agrees with the known results of this case Wannier (1950, 1973); Newell (1950); Potts (1955).
III.2.2 Imaginary-field case
@@
$Z=i^{N}B^{N}{A^{\prime}_{2}}^{N}{A^{\prime}_{3}}^{N}\sum\limits_{\left\{{\tilde{s}}\right\}=\pm 1}~{\prod\limits_{{\rm{all~vertices}}}{\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)}}$ (47)
with $A^{\prime}_{i}=\frac{1}{2}{e^{\beta J^{\prime}_{i}}}$ and
$\omega\left({{{\tilde{s}}_{1}},{{\tilde{s}}_{2}},{{\tilde{s}}_{3}},{{\tilde{s}}_{4}}}\right)=\sum\limits_{{s_{1}}=\pm 1,{s_{2}}=\pm 1}(s_{1})e^{-\beta J^{\prime}_{1}{s_{1}}{s_{2}}}\times$
$~~~~~~~~~~~~~~~~~~~~~~~~~e^{-\beta\left[{{{\tilde{J}^{\prime}}_{2}}\left({{s_{1}}{{\tilde{s}}_{2}}+{s_{2}}{{\tilde{s}}_{3}}}\right)+{{\tilde{J}^{\prime}}_{3}}\left({{s_{1}}{{\tilde{s}}_{1}}+{s_{2}}{{\tilde{s}}_{4}}}\right)}\right]}.$ (48)
@@
$\omega_{1}={e^{-\beta{J^{\prime}_{1}}}}\left[{{e^{-2\beta\left({{{\tilde{J}^{\prime}}_{2}}+{{\tilde{J}^{\prime}}_{3}}}\right)}}-{e^{2\beta\left({{{\tilde{J}^{\prime}}_{2}}+{{\tilde{J}^{\prime}}_{3}}}\right)}}}\right],$
@@
$\omega_{3}={e^{-\beta{J^{\prime}_{1}}}}\left[{{e^{2\beta\left({{{\tilde{J}^{\prime}}_{2}}-{{\tilde{J}^{\prime}}_{3}}}\right)}}-{e^{-2\beta\left({{{\tilde{J}^{\prime}}_{2}}-{{\tilde{J}^{\prime}}_{3}}}\right)}}}\right],$
$\omega_{4}=-\omega_{3},$
$\omega_{5}={e^{\beta{J^{\prime}_{1}}}}\left[{{e^{2\beta\left({{{\tilde{J}^{\prime}}_{2}}-{{\tilde{J}^{\prime}}_{3}}}\right)}}-{e^{-2\beta\left({{{\tilde{J}^{\prime}}_{2}}-{{\tilde{J}^{\prime}}_{3}}}\right)}}}\right],$
$\omega_{6}=-\omega_{5},$
$\omega_{7}={e^{\beta{J^{\prime}_{1}}}}\left[{{e^{2\beta\left({{{\tilde{J}^{\prime}}_{2}}+{{\tilde{J}^{\prime}}_{3}}}\right)}}-{e^{-2\beta\left({{{\tilde{J}^{\prime}}_{2}}+{{\tilde{J}^{\prime}}_{3}}}\right)}}}\right],$
@@
$\omega_{9}=e^{\beta J^{\prime}_{1}}\left({{e^{2\beta{\tilde{J}^{\prime}}_{2}}}-{e^{-2\beta{\tilde{J}^{\prime}}_{2}}}}\right)+e^{-\beta J^{\prime}_{1}}\left(e^{-2\beta{\tilde{J}^{\prime}}_{3}}-e^{2\beta{\tilde{J}^{\prime}}_{3}}\right),$
$\omega_{10}=-\omega_{9},$
$\omega_{11}=e^{\beta J^{\prime}_{1}}\left({{e^{2\beta{\tilde{J}^{\prime}}_{2}}}-{e^{-2\beta{\tilde{J}^{\prime}}_{2}}}}\right)+e^{-\beta J^{\prime}_{1}}\left(e^{2\beta{\tilde{J}^{\prime}}_{3}}-e^{-2\beta{\tilde{J}^{\prime}}_{3}}\right),$
$\omega_{12}=-\omega_{11},$
$\omega_{13}=e^{-\beta J^{\prime}_{1}}\left({{e^{-2\beta{\tilde{J}^{\prime}}_{2}}}-{e^{2\beta{\tilde{J}^{\prime}}_{2}}}}\right)+e^{\beta J^{\prime}_{1}}\left(e^{2\beta{\tilde{J}^{\prime}}_{3}}-e^{-2\beta{\tilde{J}^{\prime}}_{3}}\right),$
$\omega_{14}=-\omega_{13},$
$\omega_{15}=e^{-\beta J^{\prime}_{1}}\left(e^{2\beta{\tilde{J}^{\prime}}_{2}}-e^{-2\beta{\tilde{J}^{\prime}}_{2}}\right)+e^{\beta J^{\prime}_{1}}\left({{e^{2\beta{\tilde{J}^{\prime}}_{3}}}-{e^{-2\beta{\tilde{J}^{\prime}}_{3}}}}\right),$
$\omega_{16}=-\omega_{15}.$ (49)
@@
$\tilde{\omega}_{9}=2\cosh\left(\beta J^{\prime}_{1}\right)\left\{\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}+{\tilde{J}^{\prime}}_{3}\right)}\right]\right$.
$~~~~~~~~~~~~~~~~~~\left.+\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}-{\tilde{J}^{\prime}}_{3}\right)}\right]+2\sinh\left({2\beta{{\tilde{J}^{\prime}}_{2}}}\right)\right\},$
$\tilde{\omega}_{10}=2\cosh\left(\beta J^{\prime}_{1}\right)\left\{\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}+{\tilde{J}^{\prime}}_{3}\right)}\right]\right$.
$~~~~~~~~~~~~~~~~~~\left.+\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}-{\tilde{J}^{\prime}}_{3}\right)}\right]-2\sinh\left({2\beta{{\tilde{J}^{\prime}}_{2}}}\right)\right\},$
$\tilde{\omega}_{11}=-2\sinh\left(\beta J^{\prime}_{1}\right)\left\{\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}+{\tilde{J}^{\prime}}_{3}\right)}\right]\right$.
$~~~~~~~~~~~~~~~~~~\left.+\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}-{\tilde{J}^{\prime}}_{3}\right)}\right]-2\sinh\left({2\beta{{\tilde{J}^{\prime}}_{2}}}\right)\right\},$
$\tilde{\omega}_{12}=-2\sinh\left(\beta J^{\prime}_{1}\right)\left\{\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}+{\tilde{J}^{\prime}}_{3}\right)}\right]\right$.
$~~~~~~~~~~~~~~~~~~\left.+\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}-{\tilde{J}^{\prime}}_{3}\right)}\right]+2\sinh\left({2\beta{{\tilde{J}^{\prime}}_{2}}}\right)\right\},$
$\tilde{\omega}_{13}=2\cosh\left(\beta J^{\prime}_{1}\right)\left\{\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}+{\tilde{J}^{\prime}}_{3}\right)}\right]\right$.
$~~~~~~~~~~~~~~~~~~\left.-\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}-{\tilde{J}^{\prime}}_{3}\right)}\right]+2\sinh\left({2\beta{{\tilde{J}^{\prime}}_{3}}}\right)\right\},$
$\tilde{\omega}_{14}=2\cosh\left(\beta J^{\prime}_{1}\right)\left\{\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}+{\tilde{J}^{\prime}}_{3}\right)}\right]\right$.
$~~~~~~~~~~~~~~~~~~\left.-\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}-{\tilde{J}^{\prime}}_{3}\right)}\right]-2\sinh\left({2\beta{{\tilde{J}^{\prime}}_{3}}}\right)\right\},$
$\tilde{\omega}_{15}=-2\sinh\left(\beta J^{\prime}_{1}\right)\left\{\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}+{\tilde{J}^{\prime}}_{3}\right)}\right]\right$.
$~~~~~~~~~~~~~~~~~~\left.-\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}-{\tilde{J}^{\prime}}_{3}\right)}\right]-2\sinh\left({2\beta{{\tilde{J}^{\prime}}_{3}}}\right)\right\},$
$\tilde{\omega}_{16}=-2\sinh\left(\beta J^{\prime}_{1}\right)\left\{\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}+{\tilde{J}^{\prime}}_{3}\right)}\right]\right$.
$~~~~~~~~~~~~~~~~~~\left.-\sinh\left[{2\beta\left({\tilde{J}^{\prime}}_{2}-{\tilde{J}^{\prime}}_{3}\right)}\right]+2\sinh\left({2\beta{{\tilde{J}^{\prime}}_{3}}}\right)\right\}.$ (50)
It can be examined that ${\tilde{\omega}}_{9}{\tilde{\omega}}_{10}={\tilde{\omega}}_{13}{\tilde{\omega}}_{14}$ and ${\tilde{\omega}}_{11}{\tilde{\omega}}_{12}={\tilde{\omega}}_{15}{\tilde{\omega}}_{16}$, such that the odd free-fermion condition Eq. (28) is satisfied. Therefore, the triangular lattice Ising model in the imaginary field is equivalent with an odd free-fermion model. Now the result is
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z=i\frac{\pi}{2}+\ln B+\ln A^{\prime}_{2}+\ln A^{\prime}_{3}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\mathop{\lim}\limits_{{N_{{\rm{v}}}}\to\infty}\frac{1}{{{N_{{\rm{v}}}}}}\ln{Z_{{\rm{odd}}}},$ (51)
where the exact solution of the odd free-fermion model is given by Eq. (21) of Ref. Wu and Kunz (2004)
@@
$~~~~~~~~~~~\left.+c\cos\phi+d\cos\left({\theta-\phi}\right)+e\cos\left({\theta+\phi}\right)\right]$ (52)
@@
We should remark that the relation between our notation of weights and that of Ref. Wu and Kunz (2004) is
@@
Substituting $A^{\prime}_{i}=\frac{1}{2}{e^{\beta J^{\prime}_{i}}}$ and Eqs. (6), (7) and (50) into Eq. (51), we have the final expression
$\mathop{\lim}\limits_{N\to\infty}\frac{1}{N}\ln Z=i\frac{\pi}{2}+\frac{1}{{16{\pi^{2}}}}\int_{0}^{2\pi}d\theta\int_{0}^{2\pi}d\phi$
$~\ln\left[e^{-4\beta\left(J_{1}+J_{2}+J_{3}\right)}+e^{4\beta\left(J_{1}+J_{2}-J_{3}\right)}+e^{4\beta\left(J_{1}-J_{2}+J_{3}\right)}\right$.
$~~~~~~+e^{4\beta\left(-J_{1}+J_{2}+J_{3}\right)}-4-8\sinh^{2}\left({2\beta J_{3}}\right)\cos\theta$
$~~~~~~\left.+8\sinh^{2}\left({2\beta J_{2}}\right)\cos\phi-8\sinh^{2}\left({2\beta J_{1}}\right)\cos\left(\theta-\phi\right)\right].$ (54)
Then we have rederived the exact solution of this case Baxter (1965); Lin and Wu (1988), and indicated that the model in this case can be seen as an odd free-fermion model.
It is easy to find that this frustrated system (when $J_{1}=J_{2}=J_{3}>0$) in the imaginary field has a finite residual entropy in the zero temperature limit
[…]
@@
Here we set the Boltzmann constant $k_{B}=1$. Interestingly, the residual entropy in the imaginary field is half that in the zero field Wannier (1950) (we can also obtain the residual entropy in the zero field by Eq. (46), and compare the value with Eq. (55)).
@@ III.3.1 Zero-field case
$Z=\sum\limits_{\left\{{\left({{s_{1}},{s_{2}},{s_{3}},{s_{4}}}\right)}\right\}=\pm 1}~{\prod\limits_{{\rm{all~vertices}}}{\omega\left(s_{1},s_{2},s_{3},s_{4}\right)}}$ (56)
@@
$\omega\left({{s_{1}},{s_{2}},{s_{3}},{s_{4}}}\right)=\sum\limits_{{s_{5}}=\pm 1}{e^{-\beta{J_{3}}\left({{s_{1}}{s_{2}}+{s_{3}}{s_{4}}}\right)}}\times$
$~~~~~~~~~~~~~~~~~~~~~~~~~~e^{-\beta\left[{{J_{1}}{s_{5}}\left({{s_{2}}+{s_{3}}}\right)+{J_{2}}{s_{5}}\left({{s_{1}}+{s_{4}}}\right)}\right]}.$ (57)
@@
${\omega_{1}}={\omega_{2}}={e^{-2\beta{J_{3}}}}\left[{{e^{2\beta\left({{J_{1}}+{J_{2}}}\right)}}+{e^{-2\beta\left({{J_{1}}+{J_{2}}}\right)}}}\right],$
${\omega_{3}}={\omega_{4}}={e^{2\beta{J_{3}}}}\left[{{e^{2\beta\left({{J_{1}}-{J_{2}}}\right)}}+{e^{-2\beta\left({{J_{1}}-{J_{2}}}\right)}}}\right],$
${\omega_{5}}={\omega_{6}}=2{e^{2\beta J_{3}}},$
${\omega_{7}}={\omega_{8}}=2{e^{-2\beta J_{3}}},$
${\omega_{9}}=\omega_{10}=\omega_{11}=\omega_{12}={e^{2\beta J_{2}}}+{e^{-2\beta J_{2}}},$
$\omega_{13}=\omega_{14}=\omega_{15}=\omega_{16}={e^{2\beta J_{1}}}+{e^{-2\beta J_{1}}}.$ (58)
@@
${\tilde{\omega}}_{1}={e^{-2\beta{J_{3}}}}\cosh\left[{2\beta\left({{J_{1}}+{J_{2}}}\right)}\right]+{e^{2\beta{J_{3}}}}\cosh\left[{2\beta\left({{J_{1}}-{J_{2}}}\right)}\right]$
$~~~~~~~+2\cosh\left({2\beta{J_{3}}}\right)+2\cosh\left({2\beta{J_{2}}}\right)+2\cosh\left({2\beta{J_{1}}}\right),$
${\tilde{\omega}}_{2}={e^{-2\beta{J_{3}}}}\cosh\left[{2\beta\left({{J_{1}}+{J_{2}}}\right)}\right]+{e^{2\beta{J_{3}}}}\cosh\left[{2\beta\left({{J_{1}}-{J_{2}}}\right)}\right]$
$~~~~~~~+2\cosh\left({2\beta{J_{3}}}\right)-2\cosh\left({2\beta{J_{2}}}\right)-2\cosh\left({2\beta{J_{1}}}\right),$
${\tilde{\omega}}_{3}={e^{-2\beta{J_{3}}}}\cosh\left[{2\beta\left({{J_{1}}+{J_{2}}}\right)}\right]+{e^{2\beta{J_{3}}}}\cosh\left[{2\beta\left({{J_{1}}-{J_{2}}}\right)}\right]$
$~~~~~~~-2\cosh\left({2\beta{J_{3}}}\right)-2\cosh\left({2\beta{J_{2}}}\right)+2\cosh\left({2\beta{J_{1}}}\right),$
${\tilde{\omega}}_{4}={e^{-2\beta{J_{3}}}}\cosh\left[{2\beta\left({{J_{1}}+{J_{2}}}\right)}\right]+{e^{2\beta{J_{3}}}}\cosh\left[{2\beta\left({{J_{1}}-{J_{2}}}\right)}\right]$
$~~~~~~~-2\cosh\left({2\beta{J_{3}}}\right)+2\cosh\left({2\beta{J_{2}}}\right)-2\cosh\left({2\beta{J_{1}}}\right),$
${\tilde{\omega}}_{5}={\tilde{\omega}}_{6}={e^{-2\beta{J_{3}}}}\cosh\left[{2\beta\left({{J_{1}}+{J_{2}}}\right)}\right]$
$~~~~~~~~~~~~~~-{e^{2\beta{J_{3}}}}\cosh\left[{2\beta\left({{J_{1}}-{J_{2}}}\right)}\right]+2\sinh\left({2\beta{J_{3}}}\right),$
${\tilde{\omega}}_{7}={\tilde{\omega}}_{8}={e^{-2\beta{J_{3}}}}\cosh\left[{2\beta\left({{J_{1}}+{J_{2}}}\right)}\right]$
$~~~~~~~~~~~~~~-{e^{2\beta{J_{3}}}}\cosh\left[{2\beta\left({{J_{1}}-{J_{2}}}\right)}\right]-2\sinh\left({2\beta{J_{3}}}\right).$ (59)
One can examine that ${\tilde{\omega}}_{1}{\tilde{\omega}}_{2}={\tilde{\omega}}_{7}{\tilde{\omega}}_{8}$ and ${\tilde{\omega}}_{3}{\tilde{\omega}}_{4}={\tilde{\omega}}_{5}{\tilde{\omega}}_{6}$, therefore this case is an even free-fermion model. Making use of the even free-fermion solution as expressed in Eqs. (34) and (35), we give the exact result
@@
[…]
$\left.+C_{1}^{2}+C_{2}^{2}+C_{3}^{2}\right]+32S_{1}\left(S_{1}C_{2}C_{3}-C_{1}S_{2}S_{3}\right)\cos\theta$
$~~+32S_{2}\left(C_{1}S_{2}C_{3}-S_{1}C_{2}S_{3}\right)\cos\phi$
$~~\left.-32S_{3}\left(C_{1}C_{2}S_{3}-S_{1}S_{2}C_{3}\right)\cos\left(\theta-\phi\right)\right\}$
with $C_{i}=\cosh(2\beta J_{i})$ and $S_{i}=\sinh(2\beta J_{i})$ ($i=1,2,3$). This result agrees with the known expressions (see Eq. (36) of Ref. Kanô and Naya (1953) for the isotropic case for instance).
III.3.2 Imaginary-field case